OpenStudy (anonymous):

the question is.......write an equation that describes the sequence 2,7,12,17,.... then find the tenth term

myininaya (myininaya):

2=2+0(5) 7=2+1(5) 12=2+2(5) 17=2+3(5) so we have 2+i(5) if i=10, then 2+i(5)=2+10(5)=52

myininaya (myininaya):

actually that's the 11th term

myininaya (myininaya):

so you would do i=9 to find the 10th term 2+i(5)=47

OpenStudy (anonymous):

so is ur last reply the answer

myininaya (myininaya):

your questions ask for an expression to describe the sequence. do you not understand?

myininaya (myininaya):

and it asked for the tenth term if you read above you can find both

OpenStudy (anonymous):

yup thank u very much